OpenStudy (anonymous):

. Determine whether the point (1, 3) is a solution of the system 6x + 9y = 42 5x + 0y = 5 by substituting the point into each equation.

OpenStudy (amistre64):

and what do you get for each equation when you substitute the point in?

OpenStudy (anonymous):

6(1)+9(3)=42

OpenStudy (amistre64):

so far so good, does 6+27=42?

OpenStudy (anonymous):

no but the next one is 5+0=5

OpenStudy (anonymous):

you still there

OpenStudy (anonymous):

so then the answer would be no because they don't satisfy each equation

OpenStudy (anonymous):

thank you i didn't know because one of them were correct.

OpenStudy (amistre64):

if it is a solution to the system ... then it would have to make both equations true.